OpenStudy (anonymous):

Statement A-Energy is the cause and work is the effect. Statement two-Energy is a scalar but work is a vector product Options-a) Both A, B wrong b) A is right B is wrong c) A is wrong B is right d) Both A , B are right

OpenStudy (schrodinger):

Work = Force*Distance in the most generalized way possible. In terms of vectors, \[W = ||F||*||D||\cos \theta\]Where the asterisk is just plain old multiplication. \[W = F*D\]Where F and D are both vectors, and the asterisk represents a Dot Product operation. That being said, what happens whenever you take the dot product of two vectors? i.e., what is produced?

OpenStudy (anonymous):

Question is not that. Question posed by my daughter is about statement A whether statement A is Sacrosanct as internal energy in a static body does not necessarily causes work done. We need an answer which should qualify both the statements.Anyway thankx

OpenStudy (anonymous):

Thanks for your response partner. But question remains unresolved. State B that energy is a scalar and work is a vector is wrong and there is no ambiguity. In statement A it says energy is the cause and work is the effect is true but is it universally true. As my daughter puts it: Internal energy always playing havoc with the atoms and molecules but in its equilibrium state is it producing any work as we define work Force x displacement? Her argument goes beyond internal energy as she asks when she pushes the wall it does not move an inch though her body used up some energy in the process. Therefore could there be a rider in that statement A though without energy nothing could be displaced ( rather as she puts it , there has to be a transfer of energy, or conversion of one state to another state) I think we can sought help from others as well .That's Physics it is as simple and equally complicated in a developing country like India or Africa and at NASA.